# Quantifying the Effects of Parameters in Widespread SEP Events with EPREM

### v2

@@ 1 Introduction
This paper presents an analysis of the effects of multiple physical parameters in producing widespread SEP events. It uses a relatively simple model of strong magnetized shock propagating radially outward through the inner heliosphere to produce MHD quantities for a sophisticated model of proton acceleration and transport. Section 2 describes the simulation; section 3 presents the results of a baseline simulation run and seven variations, in which the value of one parameter differs from the baseline run; section 4 identifies the salient lessons to be learned by comparing the eight simulation runs; and section 5 concludes the paper.
2 EPREM
@@ 3 Simulation Results
[…]
@@
[…]
@@
[…]
@@
[…]
@@
no-diffusion 0.0 2 0.1 au 1/3 $\lambda_{\parallel}\propto|\vec{B}|^{-\beta/2}$ 100
@@ 3.1 Baseline Simulation Run
Figure 4: Fluxes of protons with energies (top-to-bottom) 1 MeV, 10 MeV, and 100 MeV at simulation days (left-to-right) 2, 3, 4, and 5. Point observer locations are indicated by black circles. Note that the streams shown are a subset of the 294 streams used in for the simulation run.
@@
Figure 3 shows simulated flux of protons with energies of 0.5, 1, 5, 10, 50, and 100 MeV at each observer during the baseline simulation run. The vertical dotted line in panels labeled 00, 01, 07, 08, 09, and 15 indicate the time at which the ideal shock passed that observer. The first point to note, from a big-picture perspective, is that substantial flux at all energies reaches most observers outside the ideal-shock cone (cf. Figure 1). Only the pair of observers separated by $180^{\circ}$ from the ideal-shock origin (panels 04 and 12) fail to record any increased flux associated with shock passage. However, both of those observers still record increased flux starting at day 4 or 5. This increased flux appears to be the result of backstreaming protons that have been accelerated by the shock on connected streams at radial distances near 2.0 au. An extended baseline simulation run (not included here), which ran to 12 days, suggests that fluxes decrease after these late post-shock increases.
@@
Figure 4 shows simulated flux of protons with 1, 10, and 100 MeV on a subset of streams at 2, 3, 4, and 5 days into the baseline simulation run. The panels of 10-MeV and 100-MeV proton flux at day 5 support the explanation that backstreaming protons contribute the flux increases at $180^{\circ}$ from the shock origin.
@@
The observers $5^{\circ}$ inside the western edge of the shock cone, record the most abrupt flux increases. After the shock passes, simulated proton fluxes broadly decay until the end of the simulation. The observer at 0.5 au records an overall decay of protons with 1-10 MeV, while protons with 50 and 100 MeV experience more of a plateau until day 4, at which point they quickly decay below observable values. The observer at 1.0 au records a flux depletion at all energies around day 4, after which all fluxes decay toward background (or at least unobservable) values. Fluxes of protons with 1-10 MeV steadily decay before and after the depletion; fluxes of protons with 50 and 100 MeV rises toward the depletion, then quickly decay afterward.
@@
The observers $5^{\circ}$ outside the western edge also record abrupt flux increases near the time of shock passage, even though the shock itself does not pass over their locations. One possible explanation could be that those observers are simply interpolating flux from nodes that are inside the shock cone, but images of the flux on nearby streams strongly suggest that this is not the case (cf. Figure 4). As the shock front propagates radially outward, it accelerates protons on streams that are well connected to those observer locations, allowing a portion of those protons to propagate sunward to the observers’ locations. Protons with 50 and 100 MeV, which require more time to acquire those energies, reach the peak values at these locations between days 3 and 4.
@@
Pairs of observers on either side of the eastern edge of the shock cone record flux profiles that are more similar than pairs of observers on either side of the western edge. As with the western-edge observers, there is a reasonable physical explanation, beyond overly aggressive interpolation, for the presence of shock-related flux increases at the locations of observers outside the shock cone. In this case, the observers are well connected to the location on the inner boundary where the shock originates, allowing protons accelerated at lower radial distances to propagate along streams to their locations. Examining the peak amplitudes of 10-MeV protons in Figure 3 and the location of increased 10-MeV flux in Figure 4 suggests that more protons propagate anti-sunward to reach the locations of observers $5^{\circ}$ outside the eastern edge of the shock cone than propagate sunward to reach the locations of observers $5^{\circ}$ outside the western edge of the shock cone.
@@
Moving farther away from the shock cone, we see greater differences both between and among pairs of observers. The pair of observers at $90^{\circ}$ from (i.e., $90^{\circ}$ east of) the shock origin exhibit increased fluxes associated with the shock passage on connected streams, but the arrival times are delayed due to propagation from the acceleration region. The 1.0-au observer is well enough connected to the shock front and the shocked region that it records extended elevated fluxes of protons with energies up to 100 MeV. The 0.5-au observer records an initial peak in protons with energies up to 50 MeV, but the high-energy protons quickly stream past and, though the shocked region is able to produce a secondary peak in protons with energies up to 10 MeV, this observer does not record significant fluxes of higher-energy protons until high-energy backstreaming protons pass its location near the end of the simulation. One final point of interest is that the peak in 1-MeV protons at 1.0 au lags the peak in higher-energy protons by at least a day. This appears to be due to a combination of delayed acceleration at nodes on the eastern edge of the shock cone, which later connect to that observer’s location, and the time necessary to propagate to that location after the streams have rotated.
@@
Fluxes recorded by the observers at $90^{\circ}$ from (i.e., $90^{\circ}$ west of) the shock origin exhibit behaviors more strongly related to the propagation of backstreaming protons. The initial spike of protons with 5 and 10 MeV recorded by the 0.5-au observer is likely due to a portion of nodes on the western edge of the shock cone briefly connecting to that location as the shock is beginning to accelerate protons to those energies. As the shock moves outward, it produces much larger fluxes of 5-MeV and 10-MeV protons, some of which stream sunward past both observer locations — first at 0.5 au, then at 1.0 au. The 0.5-au observer records a small increase in 50-MeV protons, followed by a larger increase in both 50-MeV and 100-MeV protons accelerated later in the shock’s radial evolution. The 1.0-au observer records larger corresponding increases that arrive almost a day later; both phenomena are likely due to the fact that the shock has had more time to accelerate protons to these energies and has moved out to distances better connected to that observer.
3.2 The Role of Perpendicular Diffusion
@@
These behaviors broadly apply to the point observers that straddle the eastern edge of the shock cone as well: Although the pre-shock flux profiles differ little from their baseline values, fluxes of protons with 50 and 100 MeV decay slightly faster at the 1.0-au point observers while fluxes of protons with energies $\geq 1$ MeV decay more quickly at the 0.5-au observer than in the baseline simulation run. Finally, the backstreaming flux of 50-MeV and 100-MeV protons at the very end of the simulation run seems to have disappeared with the lack of perpendicular diffusion.
@@
This trend continues to the 0.5-au observer at $90^{\circ}$, but the deviation from baseline flux profiles is significantly more drastic at the corresponding 1.0 au observer. Not only does that observer record faster post-shock decay in fluxes of protons with 50 and 100 MeV, but the flux of 5 MeV and 10 MeV protons completely drops out at approximately 3.5 days, then briefly rebounds to a fraction of its peak before falling back below observable values approximately 24 hours later. The flux of protons with $\textMeV\leq 1$$\text${MeV}} mirrors this behavior, but the resulting difference in amplitude around day 4 is not as drastic. Besides the loss of a small bump in the flux of 1-MeV protons recorded by the 0.5 au observer on day 3, the other notable effect at $90^{\circ}$ of suppressing perpendicular diffusion is again the absence of backstreaming high-energy protons during day 5.
@@
Fluxes of protons with $\textMeV\leq 10$$\text${MeV}} recorded by the point observers inside the western edge of the shock cone differ only slightly from the corresponding baseline profiles, with the exception of a sharper shock-associated peak observed at 0.5 au. The differences in fluxes of 50-MeV and 100-MeV protons are more pronounced: Both profiles exhibit much sharper initial peaks than in the baseline simulation run, and, with the exception of 50-MeV protons at 1.0 au, the depletions around day 4 are deeper. Just outside the western edge of the shock cone, the lack of perpendicular diffusion noticeably alters fluxes at both observers, with the most significant differences recorded by the observer at 0.5 au. That observer records a relatively sharp peak in fluxes of protons with $\textMeV\leq 10$$\text${MeV}} around the time when the shock passed $5^{\circ}$ east of its location, then a brief depletion that was followed by a rise to higher fluxes than recorded during the initial peak. The flux of 50-MeV protons exhibited pseudo-oscillatory behavior before rising to values consistent with the corresponding baseline post-shock profile, and the flux of 100-MeV protons appeared to exhibit similar behavior, though the initial peak either does not exist or is below observable values. The observer at 1.0 au and $5^{\circ}$ west of the shock cone recorded fluxes of protons with $\textMeV\leq 10$$\text${MeV}} that did not meaningfully differ from baseline values. However, the fluxes of 50-MeV and 100-MeV protons not only exhibit more pronounced depletions around day 3, but also rise to their initial peak approximately 6 hours later than in the baseline simulation run.
@@
Figure 6 illustrates how suppressing perpendicular diffusion removes a substantial amount of flux at all energies from streams that connect to the observers $180^{\circ}$ from the shock origin. It also suggests that perpendicular diffusion of higher-energy protons near the shock front moves them radially outward onto streams that connect with the observers $90^{\circ}$ from the shock origin, which could account for the initial spike in 50-MeV and 100-MeV proton fluxes observed in Figure 3 but not in Figure 5. Finally, it shows how perpendicular diffusion sustains fluxes, especially of lower-energy protons, on streams connected to eastern observers.
@@
3.3 The Role of Mean Free Path
@@
[…]
$\lambda_{\parallel}(|\vec{B}|)$ $=$ $\lambda_{\parallel 0}\left(\frac{\mathcal{R}}{\mathcal{R}_{0}}\right)^{\chi}\left(\frac{|\vec{B}|}{B_{0}}\right)^{-\beta/2}\quad\left(\texttt{mfpInverseB=1}\right)$
@@
To examine the role that the mean free path plays in SEP acceleration and transport, we ran five variations of the baseline simulation: one with $\textau\lambda_{\parallel 0}=1.0$$\text${au}}, one with $\chi=2/3$, one with $\beta=4.0$, one with $\beta=1.0$, and one with $\lambda(r)\propto r^{2}$.
Figure 7 shows proton fluxes from the simulation run with $\textau\lambda_{\parallel 0}=1.0$$\text${au}}. The most noticeable difference from the baseline simulation run is an overall reduction in flux at energies $\textMeV\geq 1$$\text${MeV}}, including the absence of flux at energies $\textMeV\geq 50$$\text${MeV}}. Note the shift to lower energies of fluxes shown in Figure 8. However, increasing $\lambda_{\parallel 0}$ does not simply reduce all fluxes at every observer. Most observers, especially those at 1.0 au, record earlier increases above background in fluxes at 1-10 MeV when compared to the baseline simulation run. The increased reference mean free path also leads to larger peak fluxes of protons with $\textMeV\leq 1$$\text${MeV}} observed at $90^{\circ}$, as well as an earlier peak in protons with those energies at 1.0 au. The observers at $90^{\circ}$ record fluxes of protons with 5 MeV and 10 MeV that closely resemble the 50-MeV and 100-MeV fluxes in the baseline run, suggesting that the shapes of these profiles are primarily governed by transport processes that take effect after the protons have been accelerated. Finally, all observers east of the shock origin record more backstreaming flux, albeit at lower energies, than in the baseline simulation run. Figure 8 illustrates how 10-MeV protons accelerated by the shock front near 2 au on day 4 propagate along field lines to reach the observers at $180^{\circ}$ on day 4, then the observers at $90^{\circ}$ on day 5. These effects are all consistent with an increase to $\lambda_{\parallel 0}$, which increases the mean free path across the board, allowing protons to escape from acceleration regions earlier, but with less energy, before propagating outward along magnetic field lines. It also allows accelerated protons to propagate farther in a given time frame.
@@
Figures 9 and 10 show proton fluxes from the simulation run with $\lambda\propto\mathcal{R}^{2/3}$. The overall effect of doubling the power-law dependence of mean free path on rigidity is to reduce the flux of protons with energies $\textMeV\leq 50$$\text${MeV}} and increase the flux of protons with $\textMeV\geq 50$$\text${MeV}}. The flux of 50-MeV protons appears to change very little. Two additional effects worth noting are: 1) the post-shock flux of 500-keV and 1-MeV protons recorded by the 1.0-au observers that straddle the eastern edge of the shock cone exhibits a deeper initial drop-out, compared to baseline values, before recovering to roughly baseline values; and 2) the initial increase in flux of 500-keV and 1-MeV protons recorded by the 1.0-au observer at $90^{\circ}$ from the shock origin is significantly delayed in addition to being reduced.
@@
Figures 11 and 12 show proton fluxes from the simulation run with $\beta=4.0$ (meaning that $\lambda_{\parallel}\propto|\vec{B}|^{-2}$). Overall, this change leads to larger peak fluxes of 100-MeV protons, relative to the baseline simulation run, at most point observers. This increase in high-energy flux accompanies a corresponding decrease in fluxes of protons with $\textMeV\leq 10$$\text${MeV}} at the observers straddling the western edge of the shock cone. The observers located $90^{\circ}$ from the shock origin registered the greatest overall reductions in lower-energy fluxes, while the observers located $90^{\circ}$ from the shock origin registered noticeable but less significant reductions. The observers centered on and straddling the eastern edge of the shock cone recorded similar peak values of fluxes at $\textMeV\leq 10$$\text${MeV}}, accompanied by an increase in higher-energy fluxes, in comparison to the baseline simulation run, but the post-shock falloff is much sharper. Peak fluxes at 1-10 MeV recorded by the observers straddling the western edge of the shock cone differ the least from baseline values, despite the significant increases in higher-energy fluxes, but there is a noticeably faster post-shock reduction, especially at the 0.5-au observer inside the shock cone. The observers at $180^{\circ}$ registered only modest deviations from baseline fluxes at all energies.
@@
These observations are consistent with the inverse sense of the physical reason that increasing $\beta$ when $\lambda_{\parallel}\propto|\vec{B}|^{-\beta/2}$ increased the flux of high-energy protons at the expense of lower-energy protons: A larger mean free path relative to baseline values in regions of enhanced magnetic-field strength allows protons to escape after acceleration to only a few MeV. A fraction of protons that remain in the acceleration regions for long enough to gain energies of 50 MeV or more appear with the shock passage at 1.0-au observers centered on the shock origin and straddling the eastern edge of the shock cone. Another fraction of protons accelerated to at least 50 MeV propagates sunward after acceleration beyond 1.0 au and arrives at the locations of observers west of the shock origin a few days later. The increased flux of protons with $\textMeV\leq 1$$\text${MeV}} observed east of the shock origin, where protons stream anti-sunward after acceleration at low radii, likely compensates for the decrease in higher-energy protons observed west of the shock origin.
@@
Figures 15 and 16 show proton fluxes from the simulation run with $\lambda(r)\propto r^{\beta}$ ($\beta=2$). Fluxes of protons with 1-10 MeV are effectively identical at all observer locations. Higher-energy fluxes at the $90^{\circ}$ observers are nearly identical to baseline values up to day 4, at which point the backstreaming protons reach these observers. This onset is 12-24 hours earlier than in the baseline run, depending on energy. This behavior extends to the 0.5-au observers straddling the eastern edge of the shock cone, with 1.0-au observers recording lower peak values at 50 and 100 MeV, in addition to an earlier onset of backstreaming protons. Observers west of the shock origin generally record lower fluxes at 50 and 100 MeV, with steeper post-shock decreases. The observers at $90^{\circ}$ also record slightly higher post-shock 1-MeV fluxes when compared to baseline values. The observers at $180^{\circ}$ from the shock origin record larger fluxes at energies up to 10 MeV, again corresponding to a reduction in fluxes at higher energies, with an increase at 1 MeV during day 5 that is absent in the baseline simulation run. Furthermore, the increased flux at 50 and 100 MeV peaks during day 5 and is decreasing by the end of the simulation run, while lower-energy fluxes appear to be increasing. Figure 16 confirms that, though the shock accelerates some protons to 100 MeV early in the simulation run, most of the 100-MeV flux is generated later and propagates sunward to reach the observers at $180^{\circ}$ then at $90^{\circ}$.
@@
3.4 The Role of Shock Profile

### v3

@@ I Introduction
This paper presents an analysis of the effects of multiple physical parameters in producing widespread SEP events. It uses a relatively simple model of strong magnetized shock propagating radially outward through the inner heliosphere to produce MHD quantities for a sophisticated model of proton acceleration and transport. Section II describes the simulation; section III presents the results of a baseline simulation run and seven variations, in which the value of one parameter differs from the baseline run; section IV identifies the salient lessons to be learned by comparing the eight simulation runs; and section V concludes the paper.
II EPREM
@@ III Simulation Results
All simulation runs also included 16 point observers: 8 at 0.5 au and 8 at 1.0 au. Observer azimuthal locations were chosen to cover a wide range inside and outside of the shock cone. At each radial distance, there are point observers at $0^{\circ}$, $90^{\circ}$, $180^{\circ}$, and $90^{\circ}$ from the shock origin, as well as point observers $\pm 5^{\circ}$ from each edge of the shock cone. All point-observer locations are in the solar equatorial plane. See § D for descriptions of EPREM’s stream and point observers. Figure 1 displays the radial and azimuthal locations of these point observers. Each observer’s numerical label is arbitrary but will be useful for later reference. The small yellow circle at the center represents the location of the Sun, as well as the approximate location of the inner simulation boundary. The blue arrow protruding from the Sun represents the origin of the ideal shock and the direction in which it propagates. Black dashed lines extending from the nominal solar surface indicate the region over which the outwardly propagating ideal shock extends — point observers within this region record the shocked MHD quantities while those outside do not. Finally, black dotted lines indicate the radii of point observers. Descriptions of individual simulated SEP events will at times refer to “observable” flux values. By “observable” in this context, we mean an amplitude larger than $10^{-5}$ counts $\left(\text{cm}^{2}{\text{ s}}{\text{ sr}}{\text{ MeV}}\right)^{-1}$, which we have chosen in order to be somewhat consistent with typical instrument ranges.
@@
Appendix E describes the flux spectrum used to initialize the ion distribution(s), $f_{s}$, on each simulation node. The initial spectrum thus defines the upstream, or background, population through which the ideal shock propagates. All simulation runs presented here used $J_{0}=10\ \left(\text{cm}^{2}{\text{ s}}{\text{ sr}}{\text{ MeV}}\right)^{-1}$, $r_{0}=1{\text{ au}}$, , $E_{0}=50{\text{ keV}}$, $\gamma=2$, and $E_{c}=1{\text{ MeV}}$, and exclusively modeled energetic protons.
Given the values of idealShockInitTime ($t_{s}$) and idealShockSpeed ($V_{s}$) listed in Table 1, the ideal shock reaches the radial distance of the initial outermost node, $R_{\text{max}}$, in 6 days. We therefore chose to end each simulation run after 6 days of simulated time.
@@
eMin $E_{\text{min}}$ lowest simulated ion energy 0.050 MeV
eMax $E_{\text{max}}$ highest simulated ion energy 500.0 MeV
mhdDensityAu $n_{0}$ solar-wind density at 1 au 5 $\text{cm}^{-3}$
@@
boundaryFunctAmplitude $J_{0}$ background flux of 50-keV protons 10 $\left(\text{cm}^{2}{\text{ s}}{\text{ sr}}{\text{ MeV}}\right)^{-1}$
@@
0.0 2 0.1 au 1/3 $\lambda_{\parallel}\propto|\vec{B}|^{-\beta/2}$ 100
@@ III.1 Baseline Simulation Run
Figure 4: Fluxes of protons with energies (top-to-bottom) 1 MeV, 10 MeV, and 100 MeV at simulation days (left-to-right) 2, 3, 4, and 5. Point observer locations are indicated by black circles. Note that the location of color-coded stream nodes trace out the ideal-shock cone due to the nature of the Lagrangian grid.
@@
Figure 3 shows simulated flux of protons with energies of 0.5, 1, 5, 10, 50, and 100 MeV at each observer during the baseline simulation run. The vertical dotted line in panels labeled 00, 01, 07, 08, 09, and 15 indicate the time at which the ideal shock passed that observer. The first point to note, from a big-picture perspective, is that substantial flux at all energies reaches most observers outside the ideal-shock cone (cf. Figure 1). Only the pair of observers separated by $180^{\circ}$ from the ideal-shock origin (panels 04 and 12) fail to record any increased flux associated with shock passage. However, both of those observers still record increased flux starting at day 4 or 5. This increased flux appears to be the result of. An extended baseline simulation run (not included here), which ran to 12 days, suggests that fluxes decrease after these late post-shock increases.
@@
Figure 4 shows simulated flux of protons with 1, 10, and 100 MeV on a subset of streams at 2, 3, 4, and 5 days into the baseline simulation run. The panels of 10-MeV and 100-MeV proton flux at day 5 .
@@
The observers $5^{\circ}$ inside the western edge of the shock record the most abrupt flux increases. After the shock passes, simulated proton fluxes broadly decay until the end of the simulation. The observer at 0.5 au records an overall decay of protons with 1-10 MeV, while protons with 50 and 100 MeV experience more of a plateau until day 4, at which point they quickly decay below observable values. The observer at 1.0 au records a flux depletion at all energies around day 4, after which all fluxes decay toward background (or at least unobservable) values. Fluxes of protons with 1-10 MeV steadily decay before and after the depletion; fluxes of protons with 50 and 100 MeV rises toward the depletion, then quickly decay afterward.
@@
The observers $5^{\circ}$ outside the western edge also record abrupt flux increases near the time of shock passage, even though the shock itself does not pass over their locations. One possible explanation could be that those observers are simply interpolating flux from nodes that are inside the shock cone, but images of the flux on nearby streams strongly suggest that this is not the case (cf. Figure 4).
@@
Pairs of observers on either side of the eastern edge of the shock cone record flux profiles that are more similar than pairs of observers on either side of the western edge. As with the western-edge observers, there is a reasonable physical explanation, beyond overly aggressive interpolation, for the presence of shock-related flux increases at the locations of observers outside the shock cone. In this case, the observers are well connected to the location on the inner boundary where the shock originates, allowing protons accelerated at lower radial distances to propagate along streams to their locations. Examining the peak amplitudes of 10-MeV protons in Figure 3 and the location of increased 10-MeV flux in Figure 4 suggests that more protons propagate to reach the locations of observers $5^{\circ}$ outside the eastern edge of the shock cone than to reach the locations of observers $5^{\circ}$ outside the western edge of the shock cone.
@@
Moving farther away from the shock cone, we see greater differences both between and among pairs of observers. The pair of observers at $90^{\circ}$ from (i.e., $90^{\circ}$ east of) the shock origin exhibit increased fluxes associated with the shock passage on connected streams, but the arrival times are delayed due to propagation from the acceleration region. The 1.0-au observer is well enough connected to the shock front and the shocked region that it records extended elevated fluxes of protons with energies up to MeV. The 0.5-au observer records an initial peak in protons with energies up to 50 MeV, but the high-energy protons quickly stream past and, though the shocked region is able to produce a secondary peak in protons with energies up to 10 MeV, this observer does not record significant fluxes of protons until of the simulation. One final point of interest is that the peak in 1-MeV protons at 1.0 au lags the peak in higher-energy protons by at least a day. This appears to be due to a combination of delayed acceleration at nodes on the eastern edge of the shock cone, which later connect to that observer’s location, and the time necessary to propagate to that location after the streams have rotated.
@@
Fluxes recorded by the observers at $90^{\circ}$ from (i.e., $90^{\circ}$ west of) the shock origin exhibit behaviors more strongly related to . The initial spike of protons with 5 and 10 MeV recorded by the 0.5-au observer is likely due to a portion of nodes on the western edge of the shock cone briefly connecting to that location as the shock is beginning to accelerate protons to those energies. As the shock moves outward, it produces much larger fluxes of 5-MeV and 10-MeV protons, some of which past both observer locations — first at 0.5 au, then at 1.0 au. The 0.5-au observer records a small increase in 50-MeV protons, followed by a larger increase in both 50-MeV and 100-MeV protons accelerated later in the shock’s radial evolution. The 1.0-au observer records larger corresponding increases that arrive almost a day later
III.2 The Role of Perpendicular Diffusion
@@
These behaviors broadly apply to the point observers that straddle the eastern edge of the shock cone as well: Although the pre-shock flux profiles differ little from their baseline values, fluxes of protons with 50 and 100 MeV decay slightly faster at the 1.0-au point observers while fluxes of protons with energies $\geq 1$ MeV decay more quickly at the 0.5-au observer than in the baseline simulation run. Finally, the increased flux of 50-MeV and 100-MeV protons at the very end of the simulation run seems to have disappeared with the lack of perpendicular diffusion.
@@
This trend continues to the 0.5-au observer at $90^{\circ}$, but the deviation from baseline flux profiles is significantly more drastic at the corresponding 1.0 au observer. Not only does that observer record faster post-shock decay in fluxes of protons with 50 and 100 MeV, but the flux of 5 MeV and 10 MeV protons completely drops out at approximately 3.5 days, then briefly rebounds to a fraction of its peak before falling back below observable values approximately 24 hours later. The flux of protons with $\leq 1{\text{ MeV}}$ mirrors this behavior, but the resulting difference in amplitude around day 4 is not as drastic. Besides the loss of a small bump in the flux of 1-MeV protons recorded by the 0.5 au observer on day 3, the other notable effect at $90^{\circ}$ of suppressing perpendicular diffusion is again the absence of high-energy protons during day 5.
@@
Fluxes of protons with $\leq 10{\text{ MeV}}$ recorded by the point observers inside the western edge of the shock cone differ only slightly from the corresponding baseline profiles, with the exception of a sharper shock-associated peak observed at 0.5 au. The differences in fluxes of 50-MeV and 100-MeV protons are more pronounced: Both profiles exhibit much sharper initial peaks than in the baseline simulation run, and, with the exception of 50-MeV protons at 1.0 au, the depletions around day 4 are deeper. Just outside the western edge of the shock cone, the lack of perpendicular diffusion noticeably alters fluxes at both observers, with the most significant differences recorded by the observer at 0.5 au. That observer records a relatively sharp peak in fluxes of protons with $\leq 10{\text{ MeV}}$ around the time when the shock passed $5^{\circ}$ east of its location, then a brief depletion that was followed by a rise to higher fluxes than recorded during the initial peak. The flux of 50-MeV protons exhibited pseudo-oscillatory behavior before rising to values consistent with the corresponding baseline post-shock profile, and the flux of 100-MeV protons appeared to exhibit similar behavior, though the initial peak either does not exist or is below observable values. The observer at 1.0 au and $5^{\circ}$ west of the shock cone recorded fluxes of protons with $\leq 10{\text{ MeV}}$ that did not meaningfully differ from baseline values. However, the fluxes of 50-MeV and 100-MeV protons not only exhibit more pronounced depletions around day 3, but also rise to their initial peak approximately 6 hours later than in the baseline simulation run.
@@
Figure 6 illustrates how suppressing perpendicular diffusion removes a substantial amount of flux at all energies from streams that connect to the observers $180^{\circ}$ from the shock origin. It also suggests that perpendicular diffusion of higher-energy protons. Finally, it shows how perpendicular diffusion sustains fluxes, especially of lower-energy protons, on streams connected to eastern observers.
@@
III.3 The Role of Mean Free Path
@@
$\lambda_{\parallel}(r)$ $=$ $\lambda_{\parallel 0}\left(\frac{\mathcal{R}}{\mathcal{R}_{0}}\right)^{\chi}\left(\frac{r}{\text{1 au}}\right)^{\beta}\qquad\left(\texttt{mfpInverseB=0}\right)$ (1a)
$\lambda_{\parallel}(|\vec{B}|)$ $=$ $\lambda_{\parallel 0}\left(\frac{\mathcal{R}}{\mathcal{R}_{0}}\right)^{\chi}\left(\frac{|\vec{B}|}{B_{0}}\right)^{-\beta/2}\quad\left(\texttt{mfpInverseB=1}\right)$ (1b)
@@
To examine the role that the mean free path plays in SEP acceleration and transport, we ran five variations of the baseline simulation: one with $\lambda_{\parallel 0}=1.0{\text{ au}}$, one with $\chi=2/3$, one with $\beta=4.0$, one with $\beta=1.0$, and one with $\lambda(r)\propto r^{2}$.
Figure 7 shows proton fluxes from the simulation run with $\lambda_{\parallel 0}=1.0{\text{ au}}$. The most noticeable difference from the baseline simulation run is an overall reduction in flux at energies $\geq 1{\text{ MeV}}$, including the absence of flux at energies $\geq 50{\text{ MeV}}$. Note the shift to lower energies of fluxes shown in Figure 8. However, increasing $\lambda_{\parallel 0}$ does not simply reduce all fluxes at every observer. Most observers, especially those at 1.0 au, record earlier increases above background in fluxes at 1-10 MeV when compared to the baseline simulation run. The increased reference mean free path also leads to larger peak fluxes of protons with $\leq 1{\text{ MeV}}$ observed at $90^{\circ}$, as well as an earlier peak in protons with those energies at 1.0 au. The observers at $90^{\circ}$ record fluxes of protons with 5 MeV and 10 MeV that closely resemble the 50-MeV and 100-MeV fluxes in the baseline run, suggesting that the shapes of these profiles are primarily governed by transport processes that take effect after the protons have been accelerated. Finally, all observers east of the shock origin record more, albeit at lower energies, than in the baseline simulation run.
Figure 8 illustrates how 10-MeV protons accelerated by the shock front near 2 au on day 4 propagate along field lines to reach the observers at $180^{\circ}$ on day 4, then the observers at $90^{\circ}$ on day 5. These effects are all consistent with an increase to $\lambda_{\parallel 0}$, which increases the mean free path across the board, allowing protons to escape from acceleration regions earlier, but with less energy, before propagating outward along magnetic field lines. It also allows accelerated protons to propagate farther in a given time frame.
@@
Figures 9 and 10 show proton fluxes from the simulation run with $\lambda\propto\mathcal{R}^{2/3}$. The overall effect of doubling the power-law dependence of mean free path on rigidity is to reduce the flux of protons with energies $\leq 50{\text{ MeV}}$ and increase the flux of protons with $\geq 50{\text{ MeV}}$. The flux of 50-MeV protons appears to change very little. Two additional effects worth noting are: 1) the post-shock flux of 500-keV and 1-MeV protons recorded by the 1.0-au observers that straddle the eastern edge of the shock cone exhibits a deeper initial drop-out, compared to baseline values, before recovering to roughly baseline values; and 2) the initial increase in flux of 500-keV and 1-MeV protons recorded by the 1.0-au observer at $90^{\circ}$ from the shock origin is significantly delayed in addition to being reduced.
@@
Figures 11 and 12 show proton fluxes from the simulation run with $\beta=4.0$ (meaning that $\lambda_{\parallel}\propto|\vec{B}|^{-2}$). Overall, this change leads to larger peak fluxes of 100-MeV protons, relative to the baseline simulation run, at most point observers. This increase in high-energy flux accompanies a corresponding decrease in fluxes of protons with $\leq 10{\text{ MeV}}$ at the observers straddling the western edge of the shock cone. The observers located $90^{\circ}$ from the shock origin registered the greatest overall reductions in lower-energy fluxes, while the observers located $90^{\circ}$ from the shock origin registered noticeable but less significant reductions. The observers centered on and straddling the eastern edge of the shock cone recorded similar peak values of fluxes at $\leq 10{\text{ MeV}}$, accompanied by an increase in higher-energy fluxes, in comparison to the baseline simulation run, but the post-shock falloff is much sharper. Peak fluxes at 1-10 MeV recorded by the observers straddling the western edge of the shock cone differ the least from baseline values, despite the significant increases in higher-energy fluxes, but there is a noticeably faster post-shock reduction, especially at the 0.5-au observer inside the shock cone. The observers at $180^{\circ}$ registered only modest deviations from baseline fluxes at all energies.
@@
These observations are consistent with the inverse sense of the physical reason that increasing $\beta$ when $\lambda_{\parallel}\propto|\vec{B}|^{-\beta/2}$ increased the flux of high-energy protons at the expense of lower-energy protons: A larger mean free path relative to baseline values in regions of enhanced magnetic-field strength allows protons to escape after acceleration to only a few MeV. A fraction of protons that remain in the acceleration regions for long enough to gain energies of 50 MeV or more appear with the shock passage at 1.0-au observers centered on the shock origin and straddling the eastern edge of the shock cone. Another fraction of protons accelerated to at least 50 MeV arrives at the locations of observers west of the shock origin a few days later via the combined effects of perpendicular diffusion and parallel transport along recently connected field lines. The increased flux of protons with $\leq 1{\text{ MeV}}$ observed east of the shock origin, where protons stream after acceleration at low radii, likely compensates for the decrease in higher-energy protons observed west of the shock origin..
@@
Figures 15 and 16 show proton fluxes from the simulation run with $\lambda(r)\propto r^{\beta}$ ($\beta=2$). Fluxes of protons with 1-10 MeV are effectively identical at all observer locations. Higher-energy fluxes at the $90^{\circ}$ observers are nearly identical to baseline values up to day 4, at which point protons reach these observers. This onset is 12-24 hours earlier than in the baseline run, depending on energy. This behavior extends to the 0.5-au observers straddling the eastern edge of the shock cone, with 1.0-au observers recording lower peak values at 50 and 100 MeV, in addition to an earlier onset of protons. Observers west of the shock origin generally record lower fluxes at 50 and 100 MeV, with steeper post-shock decreases. The observers at $90^{\circ}$ also record slightly higher post-shock 1-MeV fluxes when compared to baseline values. The observers at $180^{\circ}$ from the shock origin record larger fluxes at energies up to 10 MeV, again corresponding to a reduction in fluxes at higher energies, with an increase at 1 MeV during day 5 that is absent in the baseline simulation run. Furthermore, the increased flux at 50 and 100 MeV peaks during day 5 and is decreasing by the end of the simulation run, while lower-energy fluxes appear to be increasing. Figure 16 confirms that, though early in the simulation run, most of the 100-MeV flux .
@@
III.4 The Role of Shock Profile